# Distinguishing Neutron Star vs. Low-Mass Black Hole Binaries with Late Inspiral & Postmerger Gravitational Waves — Sensitivity to Transmuted Black Holes and Non-Annihilating Dark Matter

### v2

@@ 3.3 Dependence on Luminosity Distance
In Fig. 5, we present the Bayes Factor $\mathcal{B}^{\rm BNS}_{\rm LMBH}$ as a function of $D_{L}$ for the different detectors under consideration, illustrating how the discriminating power between LMBH-LMBH and BNS mergers evolves with increasing source distance. For this comparison, we present the total Bayes Factor evaluated over the full frequency range from 500 Hz to 4000 Hz. As before, we consider the BNS merger as the true signal. The strongest Bayesian evidence is observed for stiffer EoS. For the EoSs 2H & H4, the NEMO detector yields a $\mathcal{B}$ exceeding 1000 even at a distance of 200 Mpc, whereas LIGO A+ provides negligible evidence at that distance. However, with LIGO A+ and with the stiffest EoS 2H, events within 50 Mpc still yield significant evidence. For CE and ET, the evidence remains strong, exceeding a factor of 100, up to 600 Mpc for the same EoSs. The Bayesian evidence decreases progressively from stiffer to softer EoSs. For the softest EoS (2B), the evidence becomes notably small, as the NSs are sufficiently compact to closely mimic the LMBH waveform. Even with the most sensitive detectors like CE and ET, the Bayes Factor for EoS 2B is only around 11 at 200 Mpc. The intermediate EoSs exhibit evidence values that lie between those of 2H and 2B, and this trend is consistently observed across all detectors. The two finite-temperature EoSs, $\rm BHB\Lambda\phi$[97] and LS220 [98], lead to noticeable differences in the waveform already from the late inspiral phase, resulting in a huge Bayesian evidence from the inspiral phase itself. To be conservative, here we omit them.
4 Results: Constraints on Transmutation & Particle DM
4.1 Model-Independent Results on Transmutation
4.1.1 Parsing the CBC Rate into BNS & LMBH Rates
@@
As we discussed in the introduction, in current GW data analysis pipelines events with component masses in the range of $1-2.5\,M_{\odot}$ are generally categorized as BNS mergers by the LVK collaboration [35]. However, without a simultaneous EM counterpart, such as a kilonova or a short gamma-ray burst, this classification is not robust. This uncertainty becomes critical when estimating the astrophysical merger rate of BNS systems, as it opens up the possibility that some of these events may, in fact, originate from LMBH–LMBH systems. Consequently, it is important to account for this potential contamination to avoid overestimating the true BNS merger rate.
@@
The theoretical prediction of the CBC rate, based on stellar evolution and binary formation, is given as [114, 115, 116, 117],
@@
Here $t_{f}$ denotes the binary formation time, and ${d\rho_{*}}/{dt_{f}}$ denotes the cosmic star formation rate at time $t_{f}$[118]. The fraction of stellar mass that got bound as a binary is given as $\lambda\,=\,10^{-5}M_{\odot}^{-1}$[114]. The merger time delay distribution ${dP_{m}}/{dt}\propto 1/(t-t_{f})$ is the probability density that the binary formed at time $t_{f}$, will merge within a time $t-t_{f}$[114]. The time $t_{*}=4.9\times 10^{8}$ years corresponds to $z=10$, taken as the epoch for the formation of the first stars.
The above rate density can be normalized to LVK’s estimate for the same, viz., $10-1700\rm\,Gpc^{-3}\,yr^{-1}$, at the current time $t=13.8$ billion years or $z=0$. Now, as suggested earlier, this normalization is subject to the fact that there is no background in the NS-NS signal. In reality there can be LMBH-LMBH signals, which are non-distinguishable from BNSs beyond a certain $D_{L}$. If we assume there are only two candidates for a low-mass CBC, viz., NS-NS mergers & LMBH-LMBH mergers, then for a fixed $\rho_{\rm opt}$ with BNSs being the truth, the Bayes Factor $\mathcal{B}\equiv\mathcal{B}^{\rm BNS}_{\rm LMBH}=\mathcal{B}_{\rm BNS}^{\rm LMBH}$, and gives the probabilities of these CBCs either being NS-NS or LMBH-LMBH mergers as
$p({\rm BNS|data})=\frac{\mathcal{B}}{1+\mathcal{B}}\quad{\rm and}\quad p({\rm LMBH|data})=\frac{1}{1+\mathcal{B}}\,.$ (4.2)
For small values of $D_{L}$, the Bayes Factor remains sufficiently large to favor one model over the other. As $D_{L}$ increases and $\mathcal{B}$ starts decreasing, eventually tending to 1, the model probabilities approach $1/2$, indicating indistinguishability. In Fig. 6 we show the theoretical CBC rate (Eq. 4.1, solid blue line) as a function of the luminosity distance. The rate has been normalized using LVK’s current observation, which predicts the rate to be within $10-1700\,\rm Gpc^{-3}yr^{-1}$[35]. Among three different statistical models mentioned in [35], we choose the Multi-Source Model (MS), which predicts the rate to be within $130-1700\,\rm Gpc^{-3}yr^{-1}$, and for the normalization at $z=0$, we choose the mean of this model, given as $915\,\rm Gpc^{-3}yr^{-1}$.
@@
Figure 6: Merger rates as a function of $D_{L}$ observable by LIGO A+ (left panel) and ET (right panel). The blue solid line represents the theoretically predicted CBC merger rate, normalized to the latest observational limits from the LVK collaboration. The red and green dotted lines denote the most probable BNS and LMBH merger rates that would be inferred, respectively, using the Bayes Factors. Note the differing scales on the horizontal axes: with LIGO A+ reliable source classification becomes challenging beyond $\sim 120$ Mpc due to reduced SNR, while for ET this limitation occurs beyond $\sim 1200$ Mpc. For this analysis, we adopt the SLy EoS as a conservative benchmark; stiffer EoSs enhance distinguishability.
@@
Up to a source distance of approximately 400 (25) Mpc, the Bayesian evidence remains sufficiently high to reliably distinguish between BNS and LMBH mergers with ET (LIGO A+), implying that the inferred BNS merger rate is consistent with the true rate. However, beyond 400 (25) Mpc, the Bayesian evidence gradually decreases, and the probability of misidentifying LMBH-LMBH mergers as BNS mergers (or vice versa) increases. Beyond $\sim 1200$ (120) Mpc for ET (LIGO A+), the distinction becomes difficult, and the two types of sources appear nearly identical. As a result, the inferred BNS merger rate may be overestimated. As noted earlier, stiffer EoSs enhance the ability to distinguish BNS from LMBH mergers, enabling reliable separation at larger $D_{L}$ and resulting in tighter constraints.
4.1.2 Exclusion Sensitivity for Transmutation Fraction
@@
Unlike the previous section, where we estimated the rates, now we ask a different question: What is the fraction $f_{\rm LMBH}$ of the CBC rate that can be compatible with being LMBHs? To answer this question, we assume that we are given the detector specifications and the true model where all CBCs are in fact BNSs.
The expected number of CBCs ($N_{\rm CBC}$) has been computed earlier and can be expressed as [119, 114, 105],
@@
where $T$ is the years of observation, $R_{\rm CBC}$ is defined in Eq. 4.1, $D_{c}(z)$ denotes the comoving radial distance as a function of redshift, and $H(z)$ is the Hubble expansion rate at a redshift $z$. The cosmological parameters are taken from the latest Planck measurements [120]. The angular dependence of the SNR is encoded within the cumulative distribution of $\Theta$; given by $C_{\Theta}$, where the argument $\Theta\,=\,\left(\frac{\rho_{0}}{8}\frac{D_{L}(z)}{r_{0}}\right)\left(\frac{1.2\,M_{\odot}}{(1+z)\,\mathcal{M}_{c}}\right)^{5/6}$ varies within 0 and 4 [119, 121], calculated from the geometry of the sources and the antenna. $D_{L}(z)=(1+z)D_{c}(z)$ is the luminosity distance at redshift $z$, the chirp mass of the merging binary is denoted by $\mathcal{M}_{c}$, and $\rho_{0}$ and $r_{0}$ denote the SNR threshold and characteristic distance reach of a particular detector, respectively. For this analysis we have considered two cases, taking $\rho_{0}=8$ and $r_{0}=$ 80 Mpc for LIGO A+ and & 1591 Mpc for ET, respectively [121]. Apart from $R_{\rm CBC}$, the remainder of the expression in Eq. 4.3 denotes the exposure $\langle\rm VT\rangle$ for a specific detector. See Appendix A.2 for the details of the detector exposure.
@@
We perform a hypothesis test, assuming all potentially detected events will be consistent with background BNS merger events (i.e., no true LMBH merger events). In each independent redshift bin $i$, the total number of observed LMBH events is modeled as a Poisson variable with mean $\langle n_{i}\rangle=\mu S_{i}+B_{i}$, where $S_{i}$ is the expected number of LMBH (TBH in our case) merger events, and $B_{i}$ is the expected number of background BNS events which are misclassified as LMBH events, and $\mu$ is a hypothesis label with $\mu=1$ representing the presence of LMBH signals and $\mu=0$ corresponding to the null (background-only) hypothesis. We construct a log-likelihood ratio between the $\mu=1$ over $\mu=0$ hypothesis, and equate its square root to a desired statistical threshold (e.g., $Z=1.28$ for 90% confidence [122, 123], see Appendix A.3).
Figure 7: Exclusion sensitivity (90% confidence; upper limit) [122, 123] on the fraction of LMBH mergers ($f_{\rm LMBH}$) in CBCs, shown for LIGO A+ and the ET across nine luminosity distance bins, each of width 50 Mpc. The constraints improve with increasing distance due to enhanced exposure (see Fig. 11), but begin to weaken beyond $\sim 250\,\mathrm{Mpc}$ (for ET) as the distinguishability starts to decrease, reducing the ability to differentiate LMBH signals from BNS mergers, resulting in more numbers of backgrounds. LIGO A+ requires approximately 10 years of data to achieve meaningful constraints, while ET achieves stronger limits with just 1 year of observation. The exposure starts to rapidly decrease after 250 Mpc for LIGO A+, where ET can go up to very high luminosity distances.
@@
If a fraction $f_{i}$ of the CBCs are LMBHs by origin, then the signals and backgrounds for the $i^{\rm th}$ bin are given by Eq. 4.3 as,
$S_{i}=f_{i}\times R_{\text{CBC},i}\,\langle{\rm VT}\rangle_{i}\,\frac{\mathcal{B}_{i}}{1+\mathcal{B}_{i}}$
@@
where $\mathcal{B}=\mathcal{B}^{\rm LMBH}_{\rm BNS}=\mathcal{B}^{\rm BNS}_{\rm LMBH}$, assuming a fixed $\rho_{\rm opt}$. A test statistic value of $Z\geq 1.28$ corresponds to a rejection of the signal + background hypothesis over background-only hypothesis at the 90% confidence level. This threshold places a projected upper limit on the LMBH fraction, such that true values of $f_{i}$ exceeding this bound would lead to incompatibility with a BNS-only interpretation.
In Fig. 7, we present the 90% exclusion sensitivity to the LMBH merger fraction, denoted by $f_{\rm LMBH}$, as a function of $D_{L}$, for both LIGO A+ and the ET. The analysis is performed in nine bins, each with a width of 50 Mpc. The blue shaded region corresponds to 10 years of LIGO A+ observations, while the green shaded regions corresponding to ET, shown for 1 year and 10 years of observation time. For this analysis, we assume a fiducial CBC rate of $R_{\rm CBC}=1000\,\mathrm{Gpc}^{-3}\,\mathrm{yr}^{-1}$. Variations in this rate act as an overall normalization and scale the resulting constraints proportionally. As expected, the sensitivity drops at higher distances (owing to lack of distinguishability) and at very small distances (owing to small number of events). We have again adopted a conservative choice of the EoS on the softer side, specifically the SLy EoS (see Fig. 2). Choosing a stiffer EoS would lead to more pronounced matter effects in the waveform, thereby yielding comparatively stronger sensitivities.
@@ 4.2 Exclusion Sensitivity for DM-Nucleon Interaction
The preceding discussions are agnostic to the origin of LMBHs. However, now we specifically consider the scenario in which LMBHs are formed via DM capture-induced collapse of BNS systems [28, 14, 22, 31]. The LMBH merger rate becomes directly dependent on the underlying BNS merger rate, and one can obtain limits on the DM-nucleon interaction cross section $\sigma_{\chi n}$.
Non-annihilating DM particles, with mass $m_{\chi}$ and scattering cross-section with nucleons $\sigma_{\chi n}$, can be captured by a NS over its lifetime [25, 26, 31]. Considering a constant $\sigma_{\chi n}$, corresponding to a heavy mediator of the DM-nucleon interactions, the capture rate is given as [22, 32, 28],
@@
The factor involving $A^{2}={6\,m_{\chi}m_{n}}{v^{2}_{\rm{esc}}}/{\bar{v}^{2}_{\rm gal}}{(m_{\chi}-m_{n})^{2}}$ accounts for inadequate momentum transfers at larger $m_{\chi}$, given NS escape speed $v_{\rm esc}$ and typical DM density $\rho_{\chi}$ and speeds $\bar{v}_{\rm gal}$} in the galaxy. For a typical $1.35\,M_{\odot}$ NS, with 10 km radius, with a lifetime of 1 Gyr, the accumulated DM mass can reach up to $\sim 10^{-15}\,M_{\odot}$. If the captured DM particles are sufficiently massive, they can drift toward the center of the star and thermalize, forming a dense core within a thermal radius $r_{\rm th}$. For example, DM with $m_{\chi}=10^{5}$ GeV can form such a core within a radius of approximately 5 cm [28]. This overdense core may eventually undergo gravitational collapse, either due to self-gravitation or due to overcoming the fermi-degeneracy pressure, depending on whether the DM particle is bosonic or fermionic [22, 14]. The resulting collapse can lead to the formation of a small BH at the center of the NS. Once formed, the small BH at the core of the NS can eventually consume the entire star, leading to the formation of a black hole with mass comparable to that of the progenitor [30]. Such LMBHs, originating from DM capture-induced collapse of NSs, are referred to as Transmuted Black Holes (TBHs).
Figure 8: Exclusion sensitivity for DM-nucleon cross section $\sigma_{\chi n}$ as a function of DM mass $m_{\chi}$. This analysis assume bosonic non-annihilating DM, and includes all events within $D_{L}\leq 450\,\mathrm{Mpc}$. We adopt a fiducial compact binary coalescence rate of $R_{\rm CBC}=1000\,\mathrm{Gpc}^{-3}\,\mathrm{yr}^{-1}$, which acts as a normalization factor for the limits. The distinguishability between LMBH mergers and BNS mergers is computed assuming a soft nuclear EoS for the BNS mergers (SLy); adopting a stiffer EoS would yield stronger constraints due to significant mismatch between these two cases. Competing limits on this parameter space can be found in [28] where the limit was calculated assuming no background of LMBH mergers, and with considering a uniform prior on the CBC rate, which is estimated to be within $10-1700\rm Gpc^{-3}\,yr^{-1}$ till LVK’s 3rd observing run.
@@
where df/dr{df}/{dr} is the radial distribution of the progenitor BNSs in a galaxy. The transmuted black hole merger rate intrinsically depends on the BNS merger rate, as only a fraction of originally BNS systems can evolve into TBH binaries. This fraction is captured by the step function $\Theta\left[t_{0}-t_{f}-\tau_{\rm trans}(m_{\chi},\sigma_{\chi n},\rho_{\rm ext}(r,t_{0}))\right]$, which encodes the condition that the transmutation timescale $\tau_{\rm trans}=\tau_{\rm collapse}+\tau_{\rm swallow}$, which a function of the dark matter mass $m_{\chi}$, nucleon scattering cross-section $\sigma_{\chi n}$, and ambient dark matter density $\rho_{\rm ext}$, must be shorter than the time available since binary formation. Table 1 of Ref. [124] provides the brief details of the timescales required in this scenario.

### v3

@@ 2.3.2 Dependence on Luminosity Distance
In Fig. 3, we present the total Bayes factor $\mathcal{B}^{\rm BNS}_{\rm BLMBH}$ evaluated over the full frequency range from 500 Hz to 4000 Hz, as a function of $D_{L}$ for the different detectors under consideration. As before, we consider the BNS merger as the true signal. The strongest Bayesian evidence is observed for stiffer EoS. For the EoSs 2H & H4, the NEMO detector yields a $\mathcal{B}$ exceeding 1000 even at a distance of 200 Mpc, whereas LIGO A+ provides negligible evidence at that distance. However, with LIGO A+ and with the stiffest EoS 2H, events within 50 Mpc still yield significant evidence. For CE and ET, the evidence remains strong, exceeding a factor of 100, up to 600 Mpc for the same EoSs. The Bayesian evidence decreases progressively from stiffer to softer EoSs. For the softest EoS (2B), the evidence becomes notably small, as the NSs are sufficiently compact to closely mimic the LMBH waveform. Even with the most sensitive detectors like CE and ET, the Bayes Factor for EoS 2B is only around 11 at 200 Mpc. The intermediate EoSs exhibit evidence values that lie between those of 2H and 2B, and this trend is consistently observed across all detectors. The two finite-temperature EoSs, $\rm BHB\Lambda\phi$[83] and LS220 [84], lead to noticeable differences in the waveform already from the late inspiral phase, resulting in a huge Bayesian evidence from the inspiral phase itself. To be conservative, here we have omitted them.
3 Constraints on BLMBH & Particle DM
3.1 Model-Independent Results on BLMBH Abundance
3.1.1 Parsing the CBC Rate into BNS & BLMBH Rates
@@
As we discussed in the introduction, the current GW data analysis pipeline of the LVK collaboration assigns events with component masses in the range of $1-2.5\,M_{\odot}$ as BNS mergers [20]. However, without a simultaneous EM counterpart, such as a kilonova or a short gamma-ray burst, this classification is not robust. This uncertainty becomes critical when estimating the astrophysical merger rate of BNS systems, as it opens up the possibility that some of these observed low-mass events may, in fact, originate from LMBH–LMBH systems, e.g. from primordial BHs or other exotic compact objects. Consequently, it is important to account for this potential contamination to avoid overestimating the true BNS merger rate.
@@
The theoretical estimate of the CBC rate based on stellar evolution and binary formation is given as [85, 86, 87, 88]
@@
Here $t_{f}$ denotes the binary formation time, and ${d\rho_{*}}/{dt_{f}}$ denotes the cosmic star formation rate at time $t_{f}$[89]. The fraction of stellar mass that got bound as a binary is given as $\lambda\,=\,10^{-5}M_{\odot}^{-1}$[85]. The merger time delay distribution ${dP_{m}}/{dt}\propto 1/(t-t_{f})$ is the probability density that the binary formed at time $t_{f}$ will merge within a time $t-t_{f}$[85]. The time $t_{*}=4.9\times 10^{8}$ years corresponds to $z=10$, and is taken as the epoch of formation of the first stars.
@@
Figure 4: Merger rates as a function of $D_{L}$ observable by LIGO A+ (left panel) and ET (right panel). The blue solid lines show the theoretically predicted redshift dependence of the observable CBC merger rate, normalized to the latest limits from the LVK collaboration. The red and green dotted lines denote the most probable BNS and BLMBH merger rates that would be inferred, respectively, using the Bayes factors. Note the different range on the horizontal axes: with LIGO A+ reliable source classification becomes challenging beyond $\sim 120$ Mpc due to reduced SNR, while for ET this occurs beyond $\sim 1200$ Mpc. For this analysis, we adopt the SLy EoS as a conservative benchmark; stiffer EoSs enhance distinguishability.
The above rate density can be normalized to LVK’s estimate for the same, viz., $10-1700\rm\,Gpc^{-3}\,yr^{-1}$, at the current time $t=13.8$ billion years or $z=0$. Now, as remarked earlier, this normalization is subject to there being no other contribution to the CBC rate in the $1-2.5\,M_{\odot}$ except BNS mergers. In reality there can be BLMBH signals, which are non-distinguishable from BNSs beyond a certain $D_{L}$. If we assume there are only two candidates for a low-mass CBC, viz., BNS mergers & BLMBH mergers, then for a fixed $\rho_{\rm opt}$, the Bayes factor $\mathcal{B}\equiv\mathcal{B}^{\rm BNS}_{\rm BLMBH}=\mathcal{B}_{\rm BNS}^{\rm BLMBH}$. Thus, we find the probabilities of these CBCs being BNS or BLMBH mergers as
$p({\rm BNS|data})=\frac{\mathcal{B}}{1+\mathcal{B}}\quad{\rm and}\quad p({\rm BLMBH|data})=\frac{1}{1+\mathcal{B}}\,.$ (3.2)
For small values of $D_{L}$, the Bayes factor remains sufficiently large to favor one model over the other. As $D_{L}$ increases and $\mathcal{B}$ starts decreasing, eventually tending to 1, the model probabilities approach $1/2$, indicating indistinguishability. In Fig. 4 we show the theoretical CBC rate (Eq. 3.1, solid blue line) as a function of the luminosity distance. The rate has been normalized using LVK’s current estimate of the total CBC rate [20]. Of the three different statistical models mentioned in [20], here we have chosen the Multi-Source Model (MS), which predicts the rate to be within $130-1700\,\rm Gpc^{-3}yr^{-1}$, and for the normalization at $z=0$, we choose the mean of this model, given as $915\,\rm Gpc^{-3}yr^{-1}$.
@@
Up to a source distance of approximately 400 (25) Mpc, the Bayesian evidence remains sufficiently high to reliably distinguish between BNS and BLMBH mergers with ET (LIGO A+), implying that the inferred BNS merger rate is consistent with the true rate. However, beyond 400 (25) Mpc, the Bayesian evidence gradually decreases, and the probability of misidentifying BLMBH mergers as BNS mergers (or vice versa) increases. Beyond $\sim 1200$ (120) Mpc for ET (LIGO A+), the distinction becomes difficult, and the two types of sources appear nearly identical. As a result, the inferred BNS merger rate may be overestimated. As noted earlier, stiffer EoSs enhance the ability to distinguish BNS from BLMBH mergers, enabling reliable classification even at larger $D_{L}$ and resulting in tighter constraints.
3.1.2 Exclusion Sensitivity for $f_{\rm BLMBH}$
@@
Unlike the previous section, where we estimated the rates, now we ask a different question: What is the fraction $f_{\rm BLMBH}$ of the CBC rate that can be compatible with being BLMBHs? To answer this question, we assume that we are given the detector specifications and the true model where all CBCs are in fact BNSs.
The number of CBCs expected to be detected ($N_{\rm CBC}$) is [90, 85, 76],
@@
where $T$ is the years of observation, $R_{\rm CBC}$ is defined in Eq. 3.1, $D_{c}(z)$ denotes the comoving radial distance as a function of redshift, and $H(z)$ is the Hubble expansion rate at a redshift $z$. The cosmological parameters are taken from the latest Planck measurements [91]. The angular dependence of the SNR is encoded within $C_{\Theta}$ that varies within 0 and 4, calculated from the geometry of the sources and the antenna [90, 92]. Apart from $R_{\rm CBC}$, the remainder of the expression in Eq. 3.3 denotes the exposure $\langle\rm VT\rangle$ for a specific detector. See Appendix A.3 for the computation of the detector exposures for LIGO A+ and ET.
Projected Exclusion Significance
@@
We now perform a hypothesis test. In each independent redshift bin $i$, the total number of observed BLMBH events is modeled as a Poisson variable with mean $\langle n_{i}\rangle=\mu S_{i}+B_{i}$, where $S_{i}$ is the expected number of BLMBH merger events, and $B_{i}$ is the expected number of background BNS events that are misclassified as BLMBH events, and $\mu$ is a hypothesis label with $\mu=1$ representing the presence of BLMBH signals and $\mu=0$ corresponding to the null (background-only) hypothesis.
The total number of events in the $i^{\rm th}$ bin, from a Poisson process is given by
$P(n_{i})=\frac{(\mu S_{i}+B_{i})^{n_{i}}e^{-(\mu S_{i}+B_{i})}}{n_{i}!}\,$, (3.4)
where $S_{i}$ & $B_{i}$ in our case can be calculated using Eq. 3.3. The log-likelihood ratio of the signal hypothesis ($\mu=1$) over background-only hypothesis ($\mu=0$) is given by
$=-2\sum_{i}\left(B_{i}\,{\rm ln}\frac{S_{i}+B_{i}}{B_{i}}-S_{i}\right)\,.$ (3.5)
The median expected exclusion significance $Z$, i.e. the “Asimov” significance for rejecting the signal-plus-background hypothesis in against the background-only hypothesis, is given by [93, 94]
${Z}=\sqrt{q}=\sqrt{2\sum_{i}\left(B_{i}\,{\rm ln}\frac{B_{i}}{S_{i}+B_{i}}+S_{i}\right)}\,.$ (3.6)
A threshold of $Z\geq 1.28$ corresponds to a 90% confidence level. In the left panel of Fig. 5, we show how $Z$ depends on $f_{\rm BLMBH}$, when all events up to 450 Mpc are included at LIGO A+ and ET. Values of $f_{\rm BLMBH}$ larger than the shown threshold are expected to get ruled out.
Figure 5: Left panel: Exclusion significance $Z$ as a function of BLMBH fraction $f_{\rm BLMBH}$. For larger $f_{\rm BLMBH}$, the number of BLMBH-induced signal events rises relative to the background, leading to a stronger rejection of the BLMBH+BNS hypothesis over the true BNS-only hypothesis. Here we include all events up to 450 Mpc. Right panel: 90 % exclusion sensitivity to the fraction of BLMBH mergers in CBCs observable at LIGO A+ and the ET, in nine luminosity distance bins, each of width 50 Mpc. The constraints improve with increasing distance due to enhanced exposure (see Fig. 11), but begin to weaken beyond $\sim 250\,\mathrm{Mpc}$ (for ET) as the distinguishability starts to decrease, reducing the ability to reject background BNS events. The $\rm FF$ and the $\mathcal{B}$ used in these two figure panels were computed using the soft EoS SLy to be conservative; stiffer EoSs will give more promising results.
@@
If a fraction $f_{i}$ of the CBCs are LMBHs by origin, then the expected signal (BLMBH) and background (BNS) events in the $i^{\rm th}$ redshift bin are given by Eq. 3.3 as
$S_{i}=f_{i}\times R_{\text{CBC},i}\,\langle{\rm VT}\rangle_{i}\,\frac{\mathcal{B}_{i}}{1+\mathcal{B}_{i}}\,$,
@@
where $\mathcal{B}=\mathcal{B}^{\rm LMBH}_{\rm BNS}=\mathcal{B}^{\rm BNS}_{\rm LMBH}$, assuming a fixed $\rho_{\rm opt}$. We can now compute the test statistic $Z$ in each bin and obtain the expected upper limit on $f_{i}$.
In the right panel of Fig. 5, we present the 90% exclusion sensitivity to the BLMBH merger fraction, as a function of $D_{L}$, for both LIGO A+ and the ET. The analysis is performed in 9 bins, each with a width of 50 Mpc. The blue shaded region corresponds to 10 years of LIGO A+ observations, while the green shaded regions corresponding to ET, shown for 1 year and 10 years of observation time. For this analysis, we assume a fiducial CBC rate of $R_{\rm CBC}=1000\,\mathrm{Gpc}^{-3}\,\mathrm{yr}^{-1}$. This rate appears as an overall normalization and scales the resulting constraints proportionally. As expected, the sensitivity drops at higher distances (owing to lack of distinguishability) and at very small distances (owing to small number of events). We have again adopted a conservative choice of the EoS on the softer side, specifically the SLy EoS. Choosing a stiffer EoS would lead to more pronounced matter effects in the waveform, thereby yielding comparatively stronger sensitivities.
@@ 3.2 Exclusion Sensitivity for DM-Nucleon Interaction
The preceding discussions are agnostic to the origin of LMBHs. However, now we specifically consider the scenario in which LMBHs are formed via DM capture-induced collapse of BNS systems [16, 14, 95, 96]. Galactic DM with non-gravitational interactions with nucleons ($\sigma_{\chi n}$, the interaction strength) can be captured by NSs over their lifetimes [97, 98, 99, 95, 100, 101, 14, 102, 103, 104, 17, 16, 105]. Once captured, heavy DM can accumulate in the NS core, forming a dense dark core that may become gravitationally unstable and collapse into a microscopic BH. This BH can eventually consume the entire NS, ejecting approximately $10^{-3}-10^{-4}\,M_{\odot}$[106], and forming an LMBH of mass comparable to the original NS [96, 95, 100, 107, 14, 108, 109]. The BLMBH merger rate becomes directly dependent on the underlying BNS merger rate, and one can obtain limits on the DM-nucleon interaction cross section $\sigma_{\chi n}$.
Considering a constant $\sigma_{\chi n}$, corresponding to a heavy mediator of the DM-nucleon interactions, the capture rate is given as [95, 107, 16],
@@
The factor involving $A^{2}={6\,m_{\chi}m_{n}}{v^{2}_{\rm{esc}}}/{\bar{v}^{2}_{\rm gal}}{(m_{\chi}-m_{n})^{2}}$ accounts for inadequate momentum transfers at larger $m_{\chi}$, given NS escape speed $v_{\rm esc}$ and typical DM density $\rho_{\chi}$ and speeds $\bar{v}_{\rm gal}$} in the galaxy. For a typical $1.35\,M_{\odot}$ NS, with 10 km radius, with a lifetime of 1 Gyr, and with a core temperature ($T_{\rm core}$) of $2.1\times 10^{6}\,\rm K$, the accumulated DM mass can reach up to $\sim 10^{-15}\,M_{\odot}$. If the captured DM particles are sufficiently massive, they can drift toward the center of the star and thermalize, forming a dense core within a thermal radius $r_{\rm th}$. For example, DM with $m_{\chi}=10^{5}$ GeV can form such a core within a radius of approximately 5 cm [16]. This overdense core may eventually undergo gravitational collapse, either due to self-gravitation or due to overcoming the Fermi degeneracy pressure, depending on whether the DM particle is bosonic or fermionic [95, 14]. The resulting collapse can lead to the formation of a small seed BH at the center of the NS. Once formed, the seed BH at the core of the NS can eventually consume the entire star, leading to the formation of a black hole with mass comparable to that of the host star [106]. Such LMBHs, originating from DM capture-induced collapse of NSs, are referred to as transmuted black holes (TBHs).
Figure 6: Exclusion sensitivity for DM-nucleon cross section $\sigma_{\chi n}$ as a function of DM mass $m_{\chi}$. This analysis holds for non-annihilating DM (of three types: fermionic, bosonic, and bosonic with BEC formation), and includes all events within $D_{L}\leq 450\,\mathrm{Mpc}$. We adopt a fiducial merger rate of $R_{\rm CBC}=1000\,\mathrm{Gpc}^{-3}\,\mathrm{yr}^{-1}$, which acts as a normalization factor for the limits. The distinguishability between BLMBH mergers and BNS mergers is computed assuming a soft nuclear EoS for the BNS mergers (SLy); adopting a stiffer EoS would yield stronger constraints. Competing limits can be found in [16] where the limit was calculated assuming a zero background (no misclassified events) and a uniform prior on the CBC rate within $10-1700\rm Gpc^{-3}\,yr^{-1}$.
@@
where df/dr{df}/{dr} is the radial distribution of the progenitor BNSs in a galaxy. The TBH merger rate intrinsically depends on the BNS merger rate, as only a fraction of originally BNS systems can evolve into TBH binaries. This fraction is captured by the step function $\Theta\left[t_{0}-t_{f}-\tau_{\rm trans}(m_{\chi},\sigma_{\chi n},\rho_{\rm ext}(r,t_{0}))\right]$, which encodes the condition that the transmutation timescale $\tau_{\rm trans}=\tau_{\rm collapse}+\tau_{\rm swallow}$, which is a function of the dark matter mass $m_{\chi}$, nucleon scattering cross section $\sigma_{\chi n}$, and ambient dark matter density $\rho_{\rm ext}$, must be shorter than the time available since binary formation. Table 1 of Ref. [110] lists the timescales relevant to this scenario. The timescale $\tau_{\rm collapse}$ – associated with the seed BH formation from collapse of the accreted DM and depending on the nature of the DM particles: fermionic DM, bosonic DM without Bose-Einstein condensate (BEC) formation, or with BEC formation – is shown in Eq. 3.10.
$\begin{aligned} \tau_{\rm collapse}^{\rm fermion}&=1.9\times 10^{10}\,{\rm years}\left(\tfrac{10^{8}\,{\rm GeV}}{m_{\chi}}\right)\left(\tfrac{0.4\,{\rm GeV\,cm^{-3}}}{\rho_{\chi}}\right)\left(\tfrac{10^{-45}\,{\rm cm^{2}}}{\sigma_{\chi n}}\right),\\[6.0pt] \tau_{\rm collapse}^{\rm boson}&=4.8\times 10^{8}\,{\rm years}\left(\tfrac{T_{\rm core}}{2.1\times 10^{6}\,{\rm K}}\right)^{3/2}\left(\tfrac{10^{5}\,{\rm GeV}}{m_{\chi}}\right)^{3/2}\left(\tfrac{0.4\,{\rm GeV\,cm^{-3}}}{\rho_{\chi}}\right)\left(\tfrac{10^{-45}\,{\rm cm^{2}}}{\sigma_{\chi n}}\right),\\[6.0pt] \tau_{\rm collapse}^{\rm BEC}&=1.1\times 10^{9}\,{\rm years}\left(\tfrac{10^{-2}\,{\rm GeV}}{m_{\chi}}\right)^{2}\left(\tfrac{0.4\,{\rm GeV\,cm^{-3}}}{\rho_{\chi}}\right)\left(\tfrac{10^{-45}\,{\rm cm^{2}}}{\sigma_{\chi n}}\right).\end{aligned}$ (3.10)
In addition, the time taken for the seed BH to consume its host star is $\tau_{\rm swallow}\approx 3\times 10^{4}\,{\rm yr}\,(M_{\rm seed\,BH}/M_{\odot})$; this depends on the DM mass through $M_{\rm seed\,BH}$ and can exceed the $\tau_{\rm collapse}$ for heavier DM. See Ref. [16] for a detailed computation of these timescales, including possible uncertainties.